Please help me solve this problem!
Find the values of b and c in the following equation if the roots of the equation are {-6, -1}: x^2 + bx + c = 0

Answer by htmentor(1343)

Find the values of b and c in the following equation if the roots of the equation are {-6, -1}: x^2 + bx + c = 0
If the roots are x = -6 and x = -1, then we can write
(x+6)(x+1) = 0
Using FOIL, we have
x^2 + 7x + 6 = 0
So, b = 7, c = 6